# Stepwise Decomposition and Dual-stream Focus: A Novel Approach for Training-free Camouflaged Object Segmentation

### v2

@@ Abstract.
While promptable segmentation (e.g., SAM) has shown promise for various segmentation tasks, it still requires manual visual prompts for each object to be segmented. In contrast, task-generic promptable segmentation aims to reduce the need for such detailed prompts by employing only a task-generic prompt to guide segmentation across all test samples. However, when applied to Camouflaged Object Segmentation (COS), current methods still face two critical issues: 1) semantic ambiguity in getting instance-specific text prompts, which arises from insufficient discriminative cues in holistic captions, leading to foreground-background confusion; 2) semantic discrepancy combined with spatial separation in getting instance-specific visual prompts, which results from global background sampling far from object boundaries with low feature correlation, causing SAM to segment irrelevant regions. To address the issues above, we propose RDVP-MSD, a novel training-free test-time adaptation framework that synergizes Region-constrained Dual-stream Visual Prompting (RDVP) via Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT). MSD-CoT progressively disentangles image captions to eliminate semantic ambiguity, while RDVP injects spatial constraints into visual prompting and independently samples visual prompts for foreground and background points, effectively mitigating semantic discrepancy and spatial separation. Without requiring any training or supervision, RDVP-MSD achieves a state-of-the-art segmentation result on multiple COS benchmarks and delivers a faster inference speed than previous methods, demonstrating significantly improved accuracy and efficiency. The codes will be available at https://github.com/ycyinchao/RDVP-MSD
@@ 1. Introduction
Figure 1. Superior Performance and Efficiency: The proposed RDVP-MSD achieves state-of-the-art performance while requiring only about $19.85s$ per image, outperforming existing approaches (GenSAM (Hu et al., 2024a)/ProMaC (Hu et al., 2024b)) by $+7.6\%/3.7\%~{}S_{\alpha}$ and 3.6×/8.5×3.6$\times$/8.5$\times$3.6 × / 8.5 × speedup (averaged across all datasets). Results on three benchmark datasets demonstrate consistent advantages in both accuracy and computational efficiency.
Figure 2. Motivation of the proposed RDVP-MSD. (a) Text Prompts: (1) Existing methods directly extract category cues from image captions (e.g., “A camouflaged animal is hiding in the grass…”), yielding erroneous instance-specific text prompts (e.g., foreground: “grass” $\rightarrow$ misclassified). (2) Our MSD-CoT introduces a phrase-level disentanglement stage to decouple entangled semantics, purifying foreground (“snake”) and background (“grass”) prompts via MLLM-guided semantic disambiguation; (b) Visual Prompts: (1) Prior approaches depend on consistency heatmaps (foreground -- background), where globally sampled background points (blue) introduce semantic discrepancy and spatial separation, misguiding models to segment irrelevant regions. (2) The proposed RDVP independently selects high-confidence foreground (red) / background (blue) points within the object bounding box, forcing VLM to focus on discriminative regions surrounding camouflaged objects.
Existing approaches for generating VFM-compatible instance-specific visual prompts, exemplified by GenSAM (Hu et al., 2024a), employ a cyclic-generation mechanism that iteratively extracts instance-specific visual prompts through Multimodal Large Language Models (MLLMs) (Li et al., 2023a; OpenAI, 2024; Liu et al., 2023a, 2024a), coupled with Vision-Language Models (VLMs, e.g., Spatial CLIP (Hu et al., 2024a)). However, these methods exhibit fundamental limitations in complex scene understanding, particularly when target objects exhibit complete visual similarity with background textures. ProMaC (Hu et al., 2024b) alleviates this challenge by strategically leveraging hallucination priors, yet its reliance on multi-patch visual question-answering via MLLMs to filter irrelevant hallucinations introduces substantial computational overhead. As shown in Figure 1, although ProMaC achieves higher performance than GenSAM, this comes at the cost of sacrificing the efficiency of single-image inference. Our proposed method not only achieves better accuracy but also significantly improves the efficiency of single-image inference (outperforms GenSAM and ProMaC by+7.6% and+3.7% on the $S_{a}$ metric while being 3.6$\times$/8.5$\times$ faster).
@@
Contemporary task-generic promptable segmentation methods (Hu et al., 2024a, b) confront two principal problems in camouflaged scene understanding. First, existing methods that directly derive instance-specific text prompts from holistic image captions suffer from unresolved semantic ambiguity. As illustrated in Figure 2(a), camouflaged image captions (e.g., ”A camouflaged animal is hiding in the grass, blending in with the surrounding environment.”) contain insufficient discriminative cues, frequently inducing foreground-background confusion (i.e., misidentifying background elements (e.g., ”grass”) as a foreground text prompt). Our framework introduces a phase of phrase disentanglement that models contextual coexistence patterns between camouflaged objects and backgrounds. Through linguistic stepwise construction, this mechanism decomposes a holistic caption into a foreground phrase (”a small, furry creature with a mix of brown and green colors”) and a background phrase (”a grassy field with patches of brown and green grass”), followed by semantic purification via progressive MLLM interrogation to distill noise-free keywords (e.g., foreground: ”snake”). Second, previous visual prompting strategies (Hu et al., 2024a, b) based on the consistency heatmap (foreground -- background) suffer from semantic discrepancy and spatial separation, as illustrated in Figure 2(b). Specifically, global background sampling introduces semantic discrepancy (low feature correlation with the camouflaged object) and spatial separation (sampling points distant from object boundaries), causing VFMs to segment spurious regions. Our framework innovatively generates independent foreground/background heatmaps within the object bounding box, strategically selecting foreground points through corresponding instance-specific text prompts alignment maximization while sampling adversarial background points that exhibit high semantic response to background cues and spatial adjacency for focusing on the interior of the camouflaged object.
@@
In light of the issues above, we propose RDVP-MSD, a novel training-free test-time adaptation framework that synergizes Region-constrained Dual-stream Visual Prompting (RDVP) via Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT) as shown in Figure 3. The MSD-CoT mechanism implements a four-step stepwise reasoning process: 1) Caption Generation, 2) Phrase Disentanglement, 3) Keyword Identification, and 4) Coarse Location. This structured decomposition effectively resolves semantic ambiguities inherent in task-generic promptable methods, reducing the misclassification. Complementing the linguistic refinement, the Text-to-Mask Generator employs RDVP, which injects spatial constraints into visual prompting and independently samples high-confidence foreground/background points within bounding boxes. In the coarse stage, the Text-to-Mask Generator leverages phrase-level text prompts to generate coarse instance-level visual prompts via RDVP, which are fed into the VFM to produce initial segmentation masks. These masks are then refined into tighter bounding boxes. The fine-grained stage inherits word-level text prompts and the refined boxes, applying the same RDVP processing to focus on microscopic texture contrasts for pixel-accurate segmentation.
As illustrated in Table 1, leveraging MSD-CoT and RDVP, our proposed RDVP-MSD outperforms the state-of-the-art weakly supervised methods (point/scribble annotations) in Camouflaged Object Segmentation (COS). Notably, RDVP-MSD surpasses all existing task-generic promptable methods in COS benchmarks while maintaining a zero-training regime. Our principal contributions are threefold: (1) The proposed training-free test-time adaptation framework RDVP-MSD enables precise camouflaged object segmentation with faster inference speeds (8.5$\times$ faster than ProMaC); (2) Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT), which resolves semantic ambiguity in task-generic promptable methods through progressive sentence-phrase-word decomposition; (3) Region-constrained Dual-stream Visual Prompting (RDVP), which independently acquires adaptive foreground/background points within object bounding boxes, forcing VFM to focus on microscopic texture contrasts around camouflaged objects.
@@ 2.1. Camouflaged Object Segmentation
Camouflaged Object Segmentation (COS) (Fan et al., 2020) is the task of identifying and segmenting objects that exhibit high visual similarity to their background, making them difficult to identify and segment. This task is highly challenging due to the complex nature of camouflage. It has substantial practical applications in areas such as military surveillance (Truong and Tran, 2024; Wang et al., 2024b; Liu et al., 2025a), wildlife monitoring (Zhai et al., 2024; Wang et al., 2024c; Liu et al., 2025b), and autonomous driving (Zhu et al., 2025). Accurately segmenting camouflaged objects is essential for systems requiring high precision in object recognition and scene understanding.
@@ 2.2. Segment Anything Model for COS
Figure 3. Overview of the proposed RDVP-MSD framework. It comprises two core components: (1) Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT), which applies a sentence-phrase-word decomposition strategy in a four-step stepwise reasoning process: Caption Generation, Phrase Disentanglement, Keyword Identification, and Coarse Location. This process progressively refines image captions into disentangled instance-specific text prompts, resolving foreground-background ambiguity and associating bounding boxes with the corresponding semantic regions; (2) Text-to-Mask Generator, which employs Region-constrained Dual-stream Visual Prompting (RDVP) in a coarse-to-fine manner. Initially, coarse masks are generated using phrase-level text prompts, followed by refinement using discriminative word-level text prompts to achieve high-precision segmentation.
@@ 3.1. Framework Overview
As illustrated in Figure. 3, our proposed RDVP-MSD is a training-free test-time adaptation framework for segmenting camouflaged objects with only a single task-generic prompt. Specifically, given an image $X\in\mathbb{R}^{H\times W\times 3}$ containing the camouflaged scene from a test set, the RDVP-MSD generates a corresponding segmentation mask $M\in\mathbb{R}^{H\times W}$ under the task-generic prompt $P_{g}$ (e.g., camouflaged object, camouflaged animal, camouflaged entity.) setting by synergizing three frozen pre-trained models: a Multimodal Large Language Models (MLLMs, e.g., LLaVA (Liu et al., 2023a, 2024a)) for instance-specific text prompts generation, a Vision-Language Models (VLMs, e.g., Spatial CLIP (Hu et al., 2024a)) for text-to-visual prompt conversion, and a promptable Visual Foundation Models (VFMs, e.g., SAM (Kirillov et al., 2023)) for mask prediction. This process eliminates manual prompts and fine-tuning while maintaining generalization across diverse camouflaged scenarios.
@@ 3.2. Text Prompt Generation
The text prompt generation leverages the MLLM to transform a task-generic prompt $P_{g}$ into instance-specific text prompts tailored for each input image. Formally, given an input image $X\in\mathbb{R}^{H\times W\times 3}$ representing the visual content and a textual query $Q$, the MLLM with parameters $\theta$ generates a textual response $y$. The generated response $y$ explicitly characterizes instance-specific details relevant to the query.
The generation process follows an auto-regressive probabilistic modeling framework, sequentially predicting tokens based on the visual context provided by $X$, the textual query $Q$, and previously generated tokens. Mathematically, the entire response generation can be represented by the following joint probability:
(1) $p(y\mid X,Q;\theta)$ $=\prod_{t=1}^{T}p(y_{t}\mid y_{<t},X,Q;\theta),$
where $y_{t}$ denotes the token generated at time step $t$, $y_{<t}$ represents the sequence of tokens generated before time step $t$, and $T$ is the total length of the generated response. The conditional probability for each token is computed by applying a softmax function to the logits produced by the model:
(2) $y_{t}$ $\sim p(y_{t}\mid y_{<t},X,Q;\theta)$
(3) $=\text{softmax}(\text{logits}_{\theta}(y_{t}\mid y_{<t},X,Q))$
(4) $=\frac{\exp(\text{logits}_{\theta}(y_{t}\mid y_{<t},X,Q))}{\sum_{y^{\prime}}\exp(\text{logits}_{\theta}(y^{\prime}\mid y_{<t},X,Q))},$
where $\text{logits}_{\theta}(y_{t}\mid y_{<t},X,Q)$ denotes the raw predicted log-probabilities by the MLLM and $y^{\prime}$ represents all the tokens that the model could potentially predict.
Despite the advanced visual question capabilities of MLLMs, accurately generating instance-specific text prompts for camouflaged objects remains challenging due to their high visual similarity to surrounding backgrounds. Prior approaches (Hu et al., 2024a) have attempted to mitigate these challenges by incorporating image captions as textual priors. Later studies (Hu et al., 2024b) proposed using multiple local image patches to induce hallucinations in MLLMs for generating candidate knowledge, thereby reducing irrelevant hallucinations. However, as discussed in Section 1, these methods typically incur substantial computational overhead. To overcome these limitations, we introduce a novel phrase disentanglement strategy integrated into the multimodal reasoning process, termed the Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT). MSD-CoT explicitly models contextual coexistence between camouflaged objects and their backgrounds, effectively disentangling semantically entangled concepts and significantly enhancing the specificity and accuracy of the generated instance-specific text prompts.
@@ 3.2.1. Multimodal Stepwise Decomposition Chain of Thought
Chain of Thought, a method leveraging intermediate reasoning steps generated by large language models (LLMs) to enhance task-solving capabilities, has demonstrated remarkable improvements in complex NLP reasoning tasks (Kojima et al., 2022; Wei et al., 2022; Knappe et al., 2024). Recently, this paradigm has been extended to multimodal large language models (MLLMs), achieving notable performance in visual-language understanding tasks (Zhang et al., 2024a; Mitra et al., 2024; Wang et al., 2024a). However, studies (Zhang et al., 2024a) have revealed that directly generating instance-specific text prompts from image captions often leads to significant information loss, which is particularly problematic in scenarios involving highly camouflaged objects. Hence, we argue that relying solely on image captions for deriving instance-specific text prompts is suboptimal and inadequate for precise recognition of highly camouflaged objects.
@@
Due to the inherent information loss in direct image-to-caption translations (Zhang et al., 2024a), directly generating word-level text prompts from captions often introduces semantic inaccuracies, as exemplified in Figure 2(a). To overcome this limitation, we introduce an intermediate phrase disentanglement stage. This stage explicitly models the contextual coexistence between the camouflaged object and its environment, disentangling the intertwined foreground-background semantics. It generates more discriminative phrase-level text prompts:
@@
Previous studies (Hu et al., 2024b; Liu et al., 2023a, 2024a) have demonstrated that object categories generated by MLLMs can subsequently be associated with object regions via bounding box queries. We observed that for objects with lower camouflage levels, MLLMs typically produce relatively accurate bounding boxes that encompass the majority of the object. Conversely, for highly camouflaged objects, where visual features are scarcely distinguishable from the background, we assign an image-level bounding box as the coarse bounding box for the subsequent step:
@@
LLaVA1.5+SAM (Liu et al., 2024a; Kirillov et al., 2023) NeurIPS23 .662 .530 .170 .728 .501 .401 .314 .585 .666 .561 .168 .718

### v3

@@ Abstract.
While promptable segmentation (e.g., SAM) has shown promise for various segmentation tasks, it still requires manual visual prompts for each object to be segmented. In contrast, task-generic promptable segmentation aims to reduce the need for such detailed prompts by employing only a task-generic prompt to guide segmentation across all test samples. However, when applied to Camouflaged Object Segmentation (COS), current methods still face two critical issues: 1) semantic ambiguity in getting instance-specific text prompts, which arises from insufficient discriminative cues in holistic captions, leading to foreground-background confusion; 2) semantic discrepancy combined with spatial separation in getting instance-specific visual prompts, which results from global background sampling far from object boundaries with low feature correlation, causing SAM to segment irrelevant regions. To mitigate the issues above, we propose RDVP-MSD, a novel training-free test-time adaptation framework that synergizes Region-constrained Dual-stream Visual Prompting (RDVP) via Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT). MSD-CoT progressively disentangles image captions to eliminate semantic ambiguity, while RDVP injects spatial constraints into visual prompting and independently samples visual prompts for foreground and background points, effectively mitigating semantic discrepancy and spatial separation. Without requiring any training or supervision, RDVP-MSD achieves a state-of-the-art segmentation result on multiple COS benchmarks. The codes will be available at https://github.com/ycyinchao/RDVP-MSD.
Training-free Camouflaged Object Segmentation, Promptable Segmentation, Binary Segmentation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755175isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologiesccs: Computing methodologies Image segmentationccs: Computing methodologies Scene understanding
@@ 1. Introduction
Figure 1. Superior Performance and Efficiency: The proposed RDVP-MSD achieves state-of-the-art performance while requiring only about $19.85s$ per image, outperforming existing approaches (GenSAM (Hu et al., 2024a)/ProMaC (Hu et al., 2024b)) by $7.6\%/3.7\%~S_{\alpha}$ and $3.6\times/8.5\times$ speedup (averaged across all datasets). Results on three benchmark datasets demonstrate consistent advantages in both accuracy and computational efficiency.
Figure 2. Motivation of the proposed RDVP-MSD. (a) Text Prompts: (1) Existing methods directly extract category cues from image captions (e.g., “A camouflaged animal is hiding in the grass…”), yielding erroneous instance-specific text prompts (e.g., foreground: “grass” $\rightarrow$ misclassified). (2) Our MSD-CoT introduces a phrase-level disentanglement stage to decouple entangled semantics, purifying foreground (“snake”) and background (“grass”) prompts via MLLM-guided semantic disambiguation; (b) Visual Prompts: (1) Prior approaches depend on consistency heatmaps (foreground - background), where globally sampled background points (blue) introduce semantic discrepancy and spatial separation, misguiding models to segment irrelevant regions. (2) The proposed RDVP independently selects high-confidence foreground (red) / background (blue) points within the object bounding box, forcing VLM to focus on discriminative regions surrounding camouflaged objects.
@@
[…]
@@
Contemporary task-generic promptable segmentation methods (Hu et al., 2024a, b) confront two principal problems in camouflaged scene understanding. First, existing methods that directly derive instance-specific text prompts from holistic image captions suffer from unresolved semantic ambiguity. As illustrated in Figure 2(a), camouflaged image captions (e.g., ”A camouflaged animal is hiding in the grass, blending in with the surrounding environment.”) contain insufficient discriminative cues, frequently inducing foreground-background confusion (i.e., misidentifying background elements (e.g., ”grass”) as a foreground text prompt). Our framework introduces a phase of phrase disentanglement that models contextual coexistence patterns between camouflaged objects and backgrounds. Through linguistic stepwise construction, this mechanism decomposes a holistic caption into a foreground phrase (”a small, furry creature with a mix of brown and green colors”) and a background phrase (”a grassy field with patches of brown and green grass”), followed by semantic purification via progressive MLLM interrogation to distill noise-free keywords (e.g., foreground: ”snake”). Second, previous visual prompting strategies (Hu et al., 2024a, b) based on the consistency heatmap (foreground - background) suffer from semantic discrepancy and spatial separation, as illustrated in Figure 2(b). Specifically, global background sampling introduces semantic discrepancy (low feature correlation with the camouflaged object) and spatial separation (sampling points distant from object boundaries), causing VFMs to segment spurious regions. Our framework innovatively generates independent foreground/background heatmaps within the object bounding box, strategically selecting foreground points through corresponding instance-specific text prompts alignment maximization while sampling adversarial background points that exhibit high semantic response to background cues and spatial adjacency for focusing on the interior of the camouflaged object.
@@
In light of the issues above, we propose RDVP-MSD, a novel training-free test-time adaptation framework that synergizes Region-constrained Dual-stream Visual Prompting (RDVP) via Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT) as shown in Figure 3. The MSD-CoT mechanism implements a four-step stepwise reasoning process: 1) Caption Generation, 2) Phrase Disentanglement, 3) Keyword Identification, and 4) Coarse Location. This structured decomposition effectively mitigates semantic ambiguities inherent in task-generic promptable methods, reducing the misclassification. Complementing the linguistic refinement, the Text-to-Mask Generator employs RDVP, which injects spatial constraints into visual prompting and independently samples high-confidence foreground/background points within bounding boxes. In the coarse stage, the Text-to-Mask Generator leverages phrase-level text prompts to generate coarse instance-level visual prompts via RDVP, which are fed into the VFM to produce initial segmentation masks. These masks are then refined into tighter bounding boxes. The fine-grained stage inherits word-level text prompts and the refined boxes, applying the same RDVP processing to focus on microscopic texture contrasts for pixel-accurate segmentation.
[…]
@@ 2.1. Camouflaged Object Segmentation
Camouflaged Object Segmentation (COS) (Fan et al., 2020; Yin et al., 2024; Yin and Li, 2025) is the task of identifying and segmenting objects that exhibit high visual similarity to their background, making them difficult to identify and segment. This task is highly challenging due to the complex nature of camouflage. It has substantial practical applications in areas such as military surveillance (Truong and Tran, 2024; Wang et al., 2024b; Liu et al., 2025a), wildlife monitoring (Zhai et al., 2024; Wang et al., 2024c; Liu et al., 2025b), and autonomous driving (Zhu et al., 2025). Accurately segmenting camouflaged objects is essential for systems requiring high precision in object recognition and scene understanding.
@@ 2.2. Segment Anything Model for COS
Figure 3. Overview of the proposed RDVP-MSD framework. It comprises two core components: (1) Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT), which applies a sentence-phrase-word decomposition strategy in a four-step stepwise reasoning process: Caption Generation, Phrase Disentanglement, Keyword Identification, and Coarse Location. This process progressively refines image captions into disentangled instance-specific text prompts, mitigating foreground-background ambiguity and associating bounding boxes with the corresponding semantic regions. (2) Text-to-Mask Generator, which employs Region-constrained Dual-stream Visual Prompting (RDVP) in a coarse-to-fine manner. Initially, coarse masks are generated using phrase-level text prompts, followed by refinement using discriminative word-level text prompts to achieve high-precision segmentation.
@@ 3.1. Framework Overview
As illustrated in Figure. 3, our proposed RDVP-MSD is a training-free test-time adaptation framework for segmenting camouflaged objects with only a single task-generic prompt. Specifically, given an image $X\in\mathbb{R}^{H\times W\times 3}$ containing the camouflaged scene from a test set, the RDVP-MSD generates a corresponding segmentation mask $M\in\mathbb{R}^{H\times W}$ under the task-generic prompt $P_{g}$ (e.g., camouflaged object, camouflaged animal, camouflaged entity.) setting by synergizing three frozen pre-trained models: a Multimodal Large Language Models (MLLMs, e.g., LLaVA (Liu et al., 2023, 2024a)) for instance-specific text prompts generation, a Vision-Language Models (VLMs, e.g., Spatial CLIP (Hu et al., 2024a)) for text-to-visual prompt conversion, and a promptable Visual Foundation Models (VFMs, e.g., SAM (Kirillov et al., 2023)) for mask prediction. This process eliminates manual prompts and fine-tuning while maintaining generalization across diverse camouflaged scenarios.
@@ 3.2. Text Prompt Generation
The text prompt generation leverages the MLLM to transform a task-generic prompt $P_{g}$ into instance-specific text prompts tailored for each input image. Despite the advanced visual question capabilities of MLLMs, accurately generating instance-specific text prompts for camouflaged objects remains challenging due to their high visual similarity to surrounding backgrounds. Prior approaches (Hu et al., 2024a) have attempted to mitigate these challenges by incorporating image captions as textual priors. Later studies (Hu et al., 2024b) proposed using multiple local image patches to induce hallucinations in MLLMs for generating candidate knowledge, thereby reducing irrelevant hallucinations. However, as discussed in Section 1, these methods typically incur substantial computational overhead. To overcome these limitations, we introduce a novel phrase disentanglement strategy integrated into the multimodal reasoning process, termed the Multimodal Stepwise Decomposition Chain of Thought (MSD-CoT). MSD-CoT explicitly models contextual coexistence between camouflaged objects and their backgrounds, effectively disentangling semantically entangled concepts and significantly enhancing the specificity and accuracy of the generated instance-specific text prompts.
@@ 3.2.1. Multimodal Stepwise Decomposition Chain of Thought
Chain of Thought, a method leveraging intermediate reasoning steps generated by large language models (LLMs) to enhance task-solving capabilities, has demonstrated remarkable improvements in complex NLP reasoning tasks (Kojima et al., 2022; Wei et al., 2022). Recently, this paradigm has been extended to multimodal large language models (MLLMs), achieving notable performance in visual-language understanding tasks (Zhang et al., 2024a; Mitra et al., 2024; Wang et al., 2024a). However, studies (Zhang et al., 2024a) have revealed that directly generating instance-specific text prompts from image captions often leads to significant information loss, which is particularly problematic in scenarios involving highly camouflaged objects. Hence, we argue that relying solely on image captions for deriving instance-specific text prompts is suboptimal and inadequate for precise recognition of highly camouflaged objects.
@@
Due to the inherent information loss in direct image-to-caption translations (Zhang et al., 2024a), directly generating word-level text prompts from captions often introduces semantic ambiguity, as exemplified in Figure 2(a). To overcome this limitation, we introduce an intermediate phrase disentanglement stage. This stage explicitly models the contextual coexistence between the camouflaged object and its environment, disentangling the intertwined foreground-background semantics. It generates more discriminative phrase-level text prompts:
@@
Previous studies (Hu et al., 2024b; Liu et al., 2023, 2024a) have demonstrated that object categories generated by MLLMs can be associated with object regions through bounding box queries. We observed that for objects with lower camouflage levels, MLLMs typically generate relatively accurate bounding boxes that encompass the majority of the object. Conversely, for highly camouflaged objects, where visual features are difficult to distinguish from the background, we introduce a fault-tolerant mechanism. Specifically, we use an image-level bounding box as the initial coarse bounding box for providing a broader spatial constraint:
@@
LLaVA1.5+SAM (Liu et al., 2024a; Kirillov et al., 2023) CVPR24 .662 .530 .170 .728 .501 .401 .314 .585 .666 .561 .168 .718